It has been found that 1 out of every 20 people who visit a retail website make a purchase worth R500 or more. If we randomly select a sample of 15 visitors to the website, what is the probability that no more than three of the visitors will make a purchase worth R500 or more? Interpret your answer.


1
Expert's answer
2021-06-01T11:46:11-0400

Let "X=" the number of the visitors who will make a purchase worth R500 or more: "X\sim Bin(n,p)"

Given "p=1\/20=0.05, n=15."


"P(X=x)=\dbinom{n}{x}p^x(1-p)^{n-x}"

"P(X\leq 3)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)"

"\dbinom{15}{0}0.05^0(1-0.05)^{15-0}"

"+\dbinom{15}{1}0.05^1(15-0.05)^{15-1}"

"+\dbinom{15}{2}0.05^2(1-0.05)^{15-2}"

"+\dbinom{15}{3}0.05^3(15-0.05)^{15-3}"

"=0.46329123016+0.36575623434"

"+0.13475229686+0.03073297999"

"\approx0.994533"